### v1

@@ Methods
We applied this set-up on the same eight randomly chosen proteins from Nisonoff et al. [16]. For each protein, we sampled 100 sequences using either ProteinMPNN or ProteinMPNN with guidance. For each generated sequence, we define success based on two critera: (1) the generated sequence should fold into the desired structure, $\mathbf{x}$, and (2) the generated sequence should be at least as stable as the wild-type sequence, $\Delta\Delta G\geq 0$. We used AlphaFold 2 [49] to predict the structure of the generated sequences and compare the root mean square deviation (RMSD) between the predicted structure and the desired structure $\mathbf{x}$. We consider a sequence to achieve the desired fold if the RMSD is less than or equal to 22Å following previous evaluation criteria [38]. To evaluate the folding stability of our sequences, we use the $\Delta\Delta G$ regression model $p(\Delta\Delta G\,|\,\mathbf{s},\mathbf{x})$. ProteinMPNN sampling was performed with three different temperatures: $1.0$, $0.1$, and $0.01$ and we report the results for the temperature that achieved the highest success rate. For guided generation, we used a single temperature of 0.1, 200 sampling steps, sampling stochasticity of 20 and guidance factor of 100 for all proteins. On the whole, we observed that across all proteins, performing guidance with ProteinMPNN consistently improves its success rates (Fig. 2a).
Fold class-guided backbone generation
@@
First, we provide an overview of the experimental set-up for guiding ESM3 backbone structure generation with fold class labels from CATH. The purpose of this experiment to illustrate how functional annotations could be used to condition the sampling process of a zero-shot generative model without requiring modifications to the original model. We illustrate this by guiding ESM3 to sample protein backbone structures conditioned on fold class labels from CATH, a hierarchical protein structure classification scheme. This enables both high-level specification of overall secondary structure compositions and more fine-grained descriptions about particular folds. Concretely, we trained a fold classifier, $p(\mathbf{c}\,|\,\mathbf{x})$ that predict the CATH labels, $\mathbf{c}$, from the discrete structure tokens outputted by the structure encoder of ESM3, $\mathbf{x}$, on the set of CATH-labeled protein structures from the Protein Data Bank (PDB). We combined this fold class predictor with the structure generation track of ESM3, $p(\mathbf{x})$, to obtain a fold class conditioned backbone structure generative model, $p(\mathbf{x}\,|\,\mathbf{c})$.
@@
Now we describe additional details of the fold class predictor. Since ESM3 operates on discrete structure tokens, we need a predictor which predicts the fold class label of a protein from its structure token representation. To this end, we first trained a fold class predictor that takes in discrete structure tokens as inputs. The model architecture consists of an encoder-only Transformer trunk. The final per-position embedding is mean-pooled, then fed into three parallel linear layers, each projecting the embedding to a vector with the dimension equal to the number of classes for each of the three levels of CATH hierarchies: C (“Class”), A (“Architecture”) and T (“Topology/Fold”). The C level describes the overall secondary structure-content of a structure (e.g., “mainly $\alpha$”, “mainly $\beta$”, and “mixed $\alpha/\beta$), the A level roughly clusters structures by the overall arrangement of secondary structure in space (e.g.“beta barrel”, “roll”), and the T level assigns structures by how the secondary structure elements are connected and arranged (e.g.“immunoglobulin-like”, “TIM barrel”) [24]. The model was trained on the PDB dataset following the same pre-processing procedure as in Geffner et al. [29] with a total of 214,564 structures, categorized into 5 C classes, 43 A classes, and 1,336 T classes. We fed each structure through the structure encoder of ESM3 to obtain its structure token representation. The dataset is randomly divided into training, validation, and test sets at a ratio of 8:1:1, ensuring that at least one protein from each class is included in the test set whenever possible. The model is trained with a standard cross-entropy loss applied to each of the CATH level. For each protein, the loss is calculated for each level of its available CATH label. We did not systematically tune the hyper-parameters of the model. The trained fold class predictor achieves comparable accuracies on the test set as the predictor from Geffner et al. [29], which takes structure coordinates as inputs as opposed to structure tokens. The fold classifier used for guidance, $p(\mathbf{c}\,|\,\mathbf{x_{t}})$ needs to take in partially masked inputs. We initialized the model weights of this “noisy” classifier with those of the unnoised classifier, and trained it on the same training dataset as the unnoised classifier.
@@
We applied this set-up across the three hierarchies of fold class labels, following the evaluation set-up of Geffner et al. [29]. We sampled backbone structures conditioning on different classes across three levels of the CATH hierarchies. We drew 100 backbone structure samples for each C-level class, 30 samples for each A-level class, and 30 samples for each of the 30 randomly selected T-level classes, using either the unconditional structure generation track of ESM3, $p(\mathbf{x})$, or ESM3 with fold class guidance, $p(\mathbf{x}\,|\,\mathbf{c})$. For each class, for every sample, we sampled its length uniformly among the lengths present for the proteins of that class in the AFDB, with a minimum length of 50 and maximum length of 275. For unconditional generation, we used the default sampling setting of ESM3 with a temperature of 0.7 and $L$ sampling steps, where $L$ is the length of the protein sequence. We also experimented with using a lower sampling temperature and found that it results in ESM3 producing predominantly mainly $\alpha$ proteins. For guided generation, we used a temperature of 0.3, 200 sampling steps, sampling stochasticity of 20 and guidance factor of 100.
@@
We evaluated the generated structures on two criteria: (i) whether the fold classifier, $p(\mathbf{c}\,|\,\mathbf{x})$, predicts the desired class $\mathbf{c}$ as the most likely class for the generated structure $\mathbf{x}$, and (ii) whether the structures are considered designable, where designability is calculated following the procedure in Yim et al. [50]: for each generated backbone, eight sequences are sampled using ProteinMPNN [1] with a temperature of 0.1, then ESMFold [51] is used to predicted the structure of each generated sequence. The RMSD is calculated between the generated structure and the predicted structure, and a generated backbone is considered designable if the lowest RMSD—referred to as the self-consistency RMSD (scRMSD)—among the eight sampled sequences is $\leq 2$Å. For the purpose of computing aggregated metrics across multiple classes, we grouped the classes for each level by their frequency of occurrence in AFDB and calculated the reclassification accuracy and the fraction of designable proteins within each group. The frequency groups are defined in the same way as in Geffner et al. [29].

### v2

@@ Methods
We applied this set-up on the same eight randomly chosen proteins from Nisonoff et al. [16]. For each protein, we sampled 100 sequences using either ProteinMPNN or ProteinMPNN with guidance. For each generated sequence, we define success based on two criteria: (1) the generated sequence should fold into the desired structure, $\mathbf{x}$, and (2) the generated sequence should be at least as stable as the wild-type sequence, $\Delta\Delta G\leq 0$. We used AlphaFold 2 [70] to predict the structure of the generated sequences and compare the root mean square deviation (RMSD) between the predicted structure and the desired structure $\mathbf{x}$. We consider a sequence to achieve the desired fold if the RMSD is less than or equal to $2$Å following previous evaluation criteria [37]. To evaluate the folding stability of our sequences, we use the $\Delta\Delta G$ regression model $p(\Delta\Delta G\,|\,\mathbf{s},\mathbf{x})$. ProteinMPNN sampling was performed with three different temperatures: $1.0$, $0.1$, and $0.01$ and we report the results for the temperature that achieved the highest success rate. For guided generation, we used a single temperature of 0.1, Euler integration with a step size $\Delta t=0.005$, sampling stochasticity $\eta=20$ and guidance strength $\gamma=100$ for all proteins. On the whole, we observed that across all proteins, performing guidance with ProteinMPNN consistently improves its success rates (Fig. 2a). Details results are included in the Supplementary Information (Table A)
Enzyme class guided sequence generation
Overview
In this experiment, we demonstrated how guidance can generate sequences from a target enzyme classes under varying levels of sequence constraint. We trained an enzyme class predictor that classifies protein sequences into their Enzyme Commission (EC) numbers based on their sequence and used it to guide the protein language model ESM3 [4] during partial sequence redesign. EC numbers are hierarchical four-level codes describing different enzyme classes.
We use CLEAN [17] to evaluate the generated enzymes as an in silico benchmark [59]. This setup enabled us to explore the effectiveness of guidance with different sequence constraints, from complete sequence generation (no constraints) to heavily constrained redesign (50% sequence identity). We additionally present results for EC numbers of varying prevalence in the dataset, measuring generation success of the top-10 most common, 10 random, and 65 rare EC classes.
Enzyme class predictor
Our enzyme class predictor was a 3-layer multilayer perceptron (MLP) that took mean-pooled ESM3 sequence embeddings as input. The embeddings are of dimension 1536 and we used an MLP with hidden dimension ($2*1536=3072$).
The model was trained on the SwissProt dataset used by ProteInfer [71]. Unlike CLEAN [17], our model is trained on all SwissProt sequences, including non-enzymes. Sequences for each EC were randomly divided into training and test sets. During method development, we experimented with training classifiers initialized with the CLEAN weights, so the test set did not include any sequences from the CLEAN dataset. The test set contained sequences from ECs with at least 40 samples. At least 30 were used for training, and at most 10 went into the test set, as long as they weren’t in the CLEAN training data. Finally, 15% of the non-enzymes were included in the test set, and the rest were used as training data. The training data contained 315,833 samples and the testing data contained 36,354 samples. The noisy predictor’s dataset contained 10 randomly sampled noising times for each sequence.
The model was trained with a standard negative log-likelihood objective and an AdamW optimizer (learning rate of 0.001, weight decay coefficient of 0.01). We used a batch size of 15k and trained for two epochs on the full training data. The accuracy curves for each mask-level bin are plotted over the test set of the 100 most common level-4 EC numbers (Fig. 3).
Experimental setup
For each EC, we sampled 10 wild-type sequences (or as many as possible for rare ECs) from the test set and randomly masked positions according to the target masking level. Then, for each protein in our test set, we created partially masked sequences at 50-100% masked levels (50% is half the sequence fixed and 100% is generated from scratch).
In this experiment, we implemented exact classifier guidance for discrete diffusion as described in Nisonoff et al. [16] using the Gillespie sampling algorithm (Alg. 2, App. D from Peng et al. [39]). We used a guidance strength $\gamma=10$ on the classifier outputs and ESM3 sampling temperature of 0.7. We sampled all positions one at a time.
Evaluation metrics
We considered the generated sequences a success if CLEAN predicted they belonged to the target enzyme class and if they had a median pLDDT score of at least 0.8, according to the ESM3 structure prediction head.
  1. 1.
    Functional accuracy: We used CLEAN to determine if the generated sequence would be classified as belonging to the target EC class. CLEAN encodes the sequence into an embedding and is trained with a contrastive loss. It compares the embedding to a database of reference sequences from each EC class to determine a multi-class classification for each protein. It also produces probability scores with a GMM fit on their validation data. We consider the sequence to be classified correctly if the predicted class is in the set of predicted ECs from CLEAN. We have never seen a CLEAN prediction with more than three classes, although this was not investigated systematically.
  2. 2.
    Structural plausibility: We leveraged ESM3’s structure prediction head to assess whether the generated sequences would fold into stable structures. We considered a sequence structurally plausible if its predicted median pLDDT (predicted local distance difference test) score exceeded 0.80.
We analyzed each method’s performance at different masking levels (50%, 60%, 70%, 80%, 90%, and 100%) to understand how the benefit of guidance depends on the level of sequence constraint. This helped us identify regimes where guidance offered significant advantages over unguided generation. We also tested the unguided and conditional models at the 80% level for three different splits of enzyme classes (Table A). These included: the “Top-10” split of the 10 enzymes classes with the most examples in the test set, the “Random” split of 10 random enzyme classes in the test set, and the “Min-10” split consisting of the 65 most rare classes in the test set (all are among the 100 least prevalent classes in the train set and no class has $>20$ training examples.
Detailed results
The proportion of samples classified by CLEAN to be from the target (specified) enzyme class and the number of these correct samples with median pLDDT $>0.8$ were higher when using guidance. When re-designing less than 70% of the sequence, guidance has minimal effect (Fig. 2b). This aligns with the noisy predictor’s performance based on masking level (Fig. 3), where the model’s predictive accuracy plateaus by the 30% unmasking level. Based on these results, we believe guidance only prevents ESM from deviating from the target enzyme class when the sequence is sufficiently close to the target class. However, when more residues must be redesigned, guidance shows a several-fold improvement over the unguided baseline (Table A). At the 80% masking level, the Random split of 10 random target ECs shows significantly higher success rates than the Top-10 split. For the rare proteins, guidance gets higher CLEAN accuracy, but the pLDDT suffers significantly. The overall success-rate is therefore similar. Full results for each generation test split are in the Supplementary Information (Table A and A).
Fold class guided backbone generation
Overview
@@
First, we provide an overview of the experimental set-up for guiding ESM3 backbone structure generation with fold class labels from CATH [26], a hierarchical protein structure classification scheme. This enables both high-level specification of overall secondary structure compositions and more fine-grained descriptions about particular folds. Concretely, we trained a fold classifier, $p_{\phi}(\mathbf{c}\,|\,\mathbf{x})$ that predict the CATH labels, $\mathbf{c}$, from the discrete structure tokens outputted by the structure encoder of ESM3, $\mathbf{x}$, on the set of CATH-labeled protein structures from the Protein Data Bank (PDB). We combined this fold class predictor with the structure generation track of ESM3, $p_{\theta}(\mathbf{x})$, to obtain a fold class conditioned backbone structure generative model, $p_{\theta,\phi}(\mathbf{x}\,|\,\mathbf{c})$.
Fold classifier
@@
Now we describe additional details of the fold class predictor. Since ESM3 operates on discrete structure tokens, we need a predictor which predicts the fold class label of a protein from its structure token representation. To this end, we first trained a fold class predictor that takes in discrete structure tokens as inputs. The model architecture consists of an encoder-only Transformer trunk. The final per-position embedding is mean-pooled, then fed into three parallel linear layers, each projecting the embedding to a vector with the dimension equal to the number of classes for each of the three levels of CATH hierarchies: C (“Class”), A (“Architecture”) and T (“Topology/Fold”). The C level describes the overall secondary structure-content of a structure (e.g., “mainly $\alpha$”, “mainly $\beta$”, and “mixed $\alpha/\beta$), the A level roughly clusters structures by the overall arrangement of secondary structure in space (e.g.“beta barrel”, “roll”), and the T level assigns structures by how the secondary structure elements are connected and arranged (e.g.“immunoglobulin-like”, “TIM barrel”) [26]. The model was trained on the PDB dataset following the same pre-processing procedure as in Geffner et al. [18] with a total of 214,564 structures, categorized into 5 C classes, 43 A classes, and 1,336 T classes. We fed each structure through the structure encoder of ESM3 to obtain its structure token representation. The dataset is randomly divided into training, validation, and test sets at a ratio of 8:1:1, ensuring that at least one protein from each class is included in the test set whenever possible.
The model was trained with a standard cross-entropy loss applied to each of the CATH level. Some proteins have multiple CATH labels due to having multiple domains and CATH being a domain-level annotation. For such proteins, we followed Geffner et al. [18] and randomly sampled one domain label as the ground truth for each training iteration, encouraging the model to predict equal probabilities for all domain labels of a protein. During evaluation, if the model predicts any of the correct labels, the prediction is considered correct. For each protein, the loss is calculated for each level of its available CATH label. The transformer trunk of the model has a hidden dimension size of 512, 16 attention heads and 5 encoder layers. In total, the model has around 19 million parameters. The model is trained using the AdamW optimizer [72, 73] with a learning rate of 0.0001, weight decay coefficient of 0.01, a dropout rate of 0.2, a batch size of 128 and a gradient accumulation step of 4. The model is trained for around 26,000 training steps on a single NVIDIA RTX 6000 GPU. We did not systematically tune the hyperparameters of the model. On its test set, the model achieves high accuracies for each level of CATH based on the metrics of Micro Accuracy and Macro Accuracy (Table A[74].
The fold classifier used for guidance needs to take in partially masked inputs. We initialized the model weights of this “noisy” classifier with those of the unnoised classifier, and trained it on the same training dataset as the unnoised classifier. We trained the noisy classifier for around 28,000 additional training steps, with the same hyperparameters as the unnoised classifier.
Fold class guidance
@@
We performed fold class guidance with ESM3 across the three hierarchies of CATH, following the evaluation set-up of Geffner et al. [18]. We sampled backbone structures conditioning on different classes across three levels of the CATH hierarchies. We drew 100 backbone structure samples for each C-level class (mainly $\alpha$, mainly $\beta$, mixed $\alpha/\beta$), 30 samples for each A-level class, and 30 samples for each of the 30 randomly selected T-level classes, using either the unguided structure generation track of ESM3, $p_{\theta}(\mathbf{x})$, or ESM3 with fold class guidance, $p_{\theta,\phi}(\mathbf{x}\,|\,\mathbf{c})$. For each class, for every sample, we sampled its length uniformly among the lengths present for the proteins of that class in the AFDB, with a minimum length of 50 and maximum length of 275. For unguided generation, we used the default sampling setting of ESM3 with a sampling temperature of 0.7 and $L$ sampling steps, where $L$ is the length of the protein sequence. We also experimented with using a lower sampling temperature and found that it results in ESM3 producing predominantly mainly $\alpha$ proteins. For guided generation, we used a sampling temperature of 0.3, Euler integration with a step size $\Delta t=0.005$, sampling stochasticity $\eta=20$ and guidance strength $\gamma=100$. Additional results with different settings of sampling temperatures are in the Supplementary Information (Fig. 5).
@@
We evaluated the generated structures on two criteria: (i) whether the fold classifier, $p_{\phi}(\mathbf{c}\,|\,\mathbf{x})$, predicts the desired class $\mathbf{c}$ as the most likely class for the generated structure $\mathbf{x}$, and (ii) whether the structures are considered designable, where designability is calculated following the procedure in Yim et al. [75]: for each generated backbone, eight sequences are sampled using ProteinMPNN [1] with a temperature of 0.1, then ESMFold [44] is used to predicted the structure of each generated sequence. The RMSD is calculated between the generated structure and the predicted structure, and a generated backbone is considered designable if the lowest RMSD—referred to as the self-consistency RMSD (scRMSD)—among the eight sampled sequences is $\leq 2$Å. For the purpose of computing aggregated metrics across multiple classes, we grouped the classes for each level by their frequency of occurrence in AFDB [76] and calculated the classification accuracy and the fraction of designable proteins within each group (Table 5). The frequency groups are defined in the same way as in Geffner et al. [18].
Detailed results